OpenStudy (anonymous):

what is the sum of the first five terms of the geometric sequence in which a1=5 and r=1/5

OpenStudy (anonymous):

If I remember correctly, the sum of a geometric sequence is: \[\S _{n}=(a _{1}(1-r ^{n})/(1-r)\] I'm not entirely sure, though.

OpenStudy (anonymous):

express answer an improper fraction

OpenStudy (campbell_st):

Use \[s _{n} = a( 1-r^n)/(1-r)\] so \[s _{n} = 5(1-(1/5)^5)/(1-1/5)\]

OpenStudy (anonymous):

so 6.248?

OpenStudy (anonymous):

and 6248/1000 in improper fraction form

OpenStudy (anonymous):

Isn't that the formula for the nth term, rather than the one for the *sum* of the first n terms?

hero (hero):

Yes